# LDMol: Text-Conditioned Molecule Diffusion Model Leveraging Chemically Informative Latent Space

### v1

@@ Appendix A Model hyperparameters and training setup
batch size per gpu encoder: 64, decoder: 128, DiT: 16
training resources encoder: NVIDIA A100(VRAM:40GB), 15 hours decoder: NVIDIA A100(VRAM:40GB), 15 hours DiT: 8 NVIDIA A100(VRAM:40GB), 96 hours
Appendix B More detailed overview for LDMol’s application for downstream tasks
@@
Figure 5: Overall pipeline for the applications of LDMol. (a) Molecule-to-text retrieval. (b) Text-guided molecule editing. The SMILES autoencoder and the text encoder are not drawn in this figure.
@@
Figure 5-(a) shows the LDMol’s molecule-to-text retrieval process. A given query molecule is converted to a latent $z$ then a forward noise process is applied with a randomly sampled timestep $t$ and noise $\epsilon$. This $z_{t}$ is fed to LDMol with each candidate text, and the candidate that minimizes the loss $||\hat{\epsilon_{\theta}}-\epsilon||^{2}_{2}$ between $\epsilon$ and the output noise $\hat{\epsilon_{\theta}}$ is retrieved. To minimize the variance from the stochasticity of $t$ and $\epsilon$, the same process can be repeated multiple times with resampled $t$ and $\epsilon$ to use a mean loss.
@@
Figure 5-(b) illustrates the DDS-based molecule editing with LDMol. Specifically, DDS requires source data $z_{src}$, target data $z_{tgt}$ which is initialized to $z_{src}$, and their corresponding source and target text descriptions. We apply the forward noise process to $z_{src}$ and $z_{tgt}$ using the same randomly sampled $t$ and $\epsilon$ to get $z_{t,src}$ and $z_{t,tgt}$. These are fed into the pre-trained LDMol with their corresponding text, where we denote the output noise as $\hat{\epsilon}_{\theta,src}$ and $\hat{\epsilon}_{\theta,tgt}$. Finally, $z_{tgt}$ is modified towards the target text by optimizing it to the direction of $(\hat{\epsilon}_{\theta,tgt}-\hat{\epsilon}_{\theta,src})$ with a learning rate $\gamma$. Here, $\hat{\epsilon}_{\theta,tgt}$ and $\hat{\epsilon}_{\theta,tgt}$ can be replaced with the classifier-free-guidanced noises, using the output with the null text and the guidance scale $\omega$. We repeated this optimization loop 200 times with $\gamma=0.4$ and $\omega=2.0$.
Appendix C Molecule-to-text retrieval examples
@@
Figure 6: The examples of molecule-to-text retrieval result on the PCdes test set. The retrieval was done at the sentence level, and the top three retrieval outputs for each query molecule are described. The corresponding description from the data pair was correctly retrieved at first for both cases, and the other retrieved candidates weakly show a correlation with the query molecule.

### v2

@@ A.2 Model hyperparameters and training setup
batch size per GPU encoder: 64, decoder: 128, DiT: 64
A.3 More detailed procedure for LDMol’s application on downstream tasks
@@
Figure 7: Overall pipeline for the downstream task applications of LDMol. (a) Molecule-to-text retrieval. (b) Text-guided molecule editing. The SMILES autoencoder and the text encoder are not drawn in this figure.
Algorithm 1 Molecule-to-Text Retrieval with LDMol
1:$z,\mathcal{C}=\{c_{i}\}^{B}_{i=1},n\in\mathbb{N}^{+}$
2:Initialize $\text{Initialize }\texttt{Errors}[c_{i}]=0\textbf{ for }i=1\textbf{ to }B$
3:for iter=1 to $\texttt{iter}=1\textbf{ to }n$ do
4:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
5:     $z_{t}=\sqrt{\overline{\alpha}_{t}}z+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
6:     for $i=1\textbf{ to }B$ do
7:         $\texttt{Errors}[c_{i}]\mathrel{+}=||\hat{\epsilon}_{\theta}(z_{t},t,c_{i})-\epsilon||_{2}^{2}$
8:     end for
9:end for
10:Return $\text{argmin}_{c_{i}\in\mathcal{C}}\texttt{Errors}[c_{i}]$
Algorithm 2 Text-guided Molecule Editing with LDMol
1:$z_{src},c_{src},c_{tgt},N\in\mathbb{N}^{+},\gamma>0,\omega\geq 1,\mathcal{D}$
2:Initialize $\text{Initialize }z_{tgt}=z_{src}$
3:for iter=1 to $\texttt{iter}=1\textbf{ to }N$ do
4:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
5:     $z_{t,src}=\sqrt{\overline{\alpha}_{t}}z_{src}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
6:     $z_{t,tgt}=\sqrt{\overline{\alpha}_{t}}z_{tgt}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
7:     $\epsilon_{\theta,src}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,src},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,src},t,c_{src})$
8:     $\epsilon_{\theta,tgt}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,tgt},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,tgt},t,c_{tgt})$
9:     $z_{tgt}=z_{tgt}-\gamma(\epsilon_{\theta,tgt}^{\omega}-\epsilon_{\theta,src}^{\omega})$
10:end for
11:Return $\mathcal{D}(z_{tgt})$
@@
Figure 7-(a) and Algorithm 1 show the LDMol’s molecule-to-text retrieval process with a given query molecule and text candidates $\mathcal{C}=\{c_{i}\}^{B}_{i=1}$. A given query molecule is converted to a latent $z$, and then a forward noise process is applied with a randomly sampled timestep $t$ and noise $\epsilon$. This $z_{t}$ is fed to LDMol with each candidate $c_{i}$, and the candidate that minimizes the loss $||\hat{\epsilon}_{\theta}(z_{t},t,c_{i})-\epsilon||^{2}_{2}$ between $\epsilon$ and the output noise $\hat{\epsilon}_{\theta}(z_{t},t,c_{i})$ is retrieved. To minimize the variance from the stochasticity of $t$ and $\epsilon$, the same process can be repeated $n$ times with resampled $t$ and $\epsilon$ to use a mean loss.
@@
Figure 7-(b) and Algorithm 2 illustrate the DDS-based molecule editing with LDMol. Specifically, DDS requires source data $z_{src}$, target data $z_{tgt}$ which is initialized to $z_{src}$, and their corresponding source and target text descriptions $\{c_{src},c_{tgt}\}$. We apply the forward noise process to $z_{src}$ and $z_{tgt}$ using the same randomly sampled $t$ and $\epsilon$ to get $z_{t,src}$ and $z_{t,tgt}$. These are fed into the pre-trained LDMol with their corresponding text, where we denote the output noise as $\hat{\epsilon}_{\theta,src}$ and $\hat{\epsilon}_{\theta,tgt}$. Finally, $z_{tgt}$ is modified towards the target text by optimizing it to the direction of $(\hat{\epsilon}_{\theta,tgt}-\hat{\epsilon}_{\theta,src})$ with a learning rate $\gamma$. Here, $\hat{\epsilon}_{\theta,tgt}$ and $\hat{\epsilon}_{\theta,tgt}$ can be replaced with the classifier-free-guidanced noises, utilizing the output with the null text and the guidance scale $\omega$. $z_{tgt}$ is decoded back as the editing output after the optimization step is iterated $N$ times. In Figure 5-(a), where we applied the same scenario to a batch of molecules, we used null text as $c_{src}$ since it’s impractical to prepare a source prompt for each molecule. The hyperparameters $\{N,\gamma,\omega\}$ are fixed for each scenario, where every choice is in the range of $100\leq N\leq 200$, $0.1\leq\gamma\leq 0.3$ and $2.0\leq\omega\leq 4.5$. For the illustrated examples in Figure 5-(b), we repeated DDS iterations with $N=150$, $\gamma=0.1$ and $\omega=2.5$.
Appendix B Additional results
B.1 Effect of classifier-free guidance scale on LDMol sample quality
Figure 8 plots the LDMol’s text-to-molecule generation performance on the ChEBI-20 test set, with different classifier-free guidance scale $\omega$ in the sampling process. Starting from $\omega=1.0$ which is equivalent to a naive conditional generation, we observed the overall sample quality is improved as $\omega$ increases but collapses for too big $\omega$. This agrees with the well-known observation on image diffusion models, and we decided to use $\omega=2.5$ for the text-to-molecule generation with LDMol.
Figure 8: Text-to-molecule generation performance of LDMol against different classifier-free guidance scales.
B.2 Molecule-to-text retrieval examples
@@
Figure 9 contains examples of molecule-to-text retrieval results with molecules from the PCdes test set. The retrieval was done at the sentence level, and the top three retrieval outputs for each query molecule are described. The corresponding description from the data pair was correctly retrieved at first for all cases, and the other retrieved candidates show a weak correlation with the query molecule.
Figure 9: The examples of molecule-to-text retrieval result on the PCdes test set. Three sentences with the lowest noise estimation error were retrieved for each query molecule.